4. What is the pH of 2.00mL of 1.0 M HCI solution diluted to 50.0 mL? ( 2pts)
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 1 M

V1 = 2 mL

V2 = 50 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (1*2)/50

M2 = 0.040 M

use:

pH = -log [H+]

= -log (4*10^-2)

= 1.3979

Answer: 1.40